HOW (1-q1/q2) expression for effeciency in carnot engine became(1-T1/T2)

HOW  (1-q1/q2) expression for effeciency in carnot engine became(1-T1/T2)

Grade:12

2 Answers

Saurabh Koranglekar
3 years ago
Dear student

Please find the link below

https://www.askiitians.com/revision-notes/physics/thermodynamics/

Regards
Vikas TU
3 years ago
Dear student 
Q2/Q1 = KT2Log(V3/V4) / KT1Log(V2/V1)
So Q2/Q1 = T2/T1 
So efficiency = 1-T2/T1
Good Luck 
Cheers